OpenStudy (anonymous):

change -3x+2y=8 into slope intercept form.be sure to show all work,and i dentify the slope and y intercept of the equation.

OpenStudy (aravindg):

\[\huge \text{WELCOME TO OPENSTUDY!}\]

OpenStudy (anonymous):

y = mx +c m = slope c= y intercept

OpenStudy (anonymous):

ax+by = c y = -a/bx + c/b ----> slope intercept form here a = -3 , b= 2 and c = 8 do it now!!!!

OpenStudy (anonymous):

-3x+2y=8 2y = 3x + 8 y = 3x/2 + 4

OpenStudy (anonymous):

Hope nw u can Solve it